# Stop-and-go waves reconstruction via iterative refinement

### v1

@@ 4.1 I-24 MOTION data
Trajectory data produced by I-24 MOTION was used to create fine-grained mean speed field data for this work. A selection of 33 days were selected, recorded from May 17 to July 17, 2024; for each day the morning rush hour period (6:00 or 5:30 AM to 10:00 AM) was considered as this period has the most dynamic and variable traffic conditions. Figure 3 shows the raw trajectory data for each selected day. The selected trajectory data was then sampled to create a fine-grained mean speed field with 4 second and 0.02 mile bin resolution. For cells with no data and cells with inconsistent values, speed was interpolated using an adaptive smoothing method [43; 28], with parameters shown in Table 2.
@@
Table 1 provides an overview of the key metrics in the I-24 MOTION dataset, including Vehicle Miles Traveled (VMT), Vehicle Hours Traveled (VHT), and the mean speed (MS, the vehicle-miles-weighted speed) [63]. It also outlines the strategy for splitting the data into training and validation sets, with a 75% allocation for training and 25% for validation. Note that the testing set comprises data from 5 days with distinctive congestion patterns due to observable events, which are excluded from both the training and validation sets to avoid model training imbalance.
@@ 4.2 I-24 RDS data
The I-24 RDS data is obtained from the I-24 radar detection system (RDS) [64; 65]. The system collects lane-level local traffic measurements including speed, occupancy, and volume, aggregated at 30-second intervals. All the radar sensors utilized in this study were meticulously calibrated. For the purposes of this study, we specifically selected the subset of sensors that align spatially with the I-24 MOTION testbed as shown in Figure 4. A total of 11 RDS sensors are aligned in this section. Two additional sensors located at both ends are selected to help with reconstruction.
@@ 4.3 Dataset processing
$\sigma$(mile) smoothing width in time coordinate 0.12 1
$\tau$(second) smoothing width in space coordinate 20 60
$c_{\text{free}}$(mph) wave speed in free traffic -12.5 -13.0
$c_{\text{cong}}$(mph) wave speed in congested traffic 60.0 60.0
$V_{\text{thr}}$(mph) crossover from congested to free traffic 37.29 37.29
$\Delta V$(mph) transition width between congested and free traffic 12.43 12.43
Dataset augmentation. To augment the dataset, we employed a sliding window strategy, generating multiple overlapping sub-datasets from the original daily dataset matrix. Starting with an original tensor of size 3600 $\times$ 200, we created multiple 200 $\times$ 200 sub-tensors. A 200 $\times$ 200 window slides across the larger matrix with a defined step size, producing a new 200 $\times$ 200 tensor at each step. In this study, we used a sliding window step size of 10, corresponding to 40 seconds in time. For the validation part, the augmentation via sliding window is not used.
@@ 4.4 Data Availability
All days of I-24 MOTION mean speed field data, as well as all corresponding days of I-24 RDS data, will be made available at i24motion.org.
@@ 5.1 Task specific architectural details
In our problem, the coarse-grained data $\mathbf{r}$ is preprocessed with upsampling to be $\mathbf{\tilde{r}}$ (as shown in Figure 5) is concatenated with the noisy fine-grained speed data $\mathbf{m}_{t}$. This process generates a 4-channel tensor as the model input, while the model output is a single-channel tensor $\mathbf{m}_{t-1}$. The UNet [61] architecture employed in this paper comprises an initial level with 64 channels, followed by levels with 128, 256, and 512 channels, culminating in a bottleneck level with 1024 channels, as shown in Figure 5. Each block within this architecture consists of three convolutional ResNet blocks augmented with a multi-head attention mechanism, leading to a total of 130 million trainable parameters.
@@
In this paper, we define a sequence of $\gamma$ values, which are essential for our diffusion process. These values are uniformly distributed with $p(\gamma)$ within a specified range for $T$ steps, starting from $\gamma_{\text{start}}=0.001$ to $\gamma_{\text{end}}=0.02$. The total number of values in this sequence is equal to the number of time steps, $T=500$, in our diffusion process. The learning rate for the optimizer of the neural denoising model is set at $2\times 10^{-4}$. The velocity $v_{c}$ is set to 16 mph, and the parameter $\omega$ is set to 0.8. An early stopping strategy is employed when the validation loss begins to increase. The best performance on the validation dataset is retained for inference purposes.
@@ 5.2 Benchmarked methods
  1. (1)
    Average: missing data points are imputed as the average of all available data points, assuming the data is uniformly distributed in space and time.
  2. (2)
@@
  4. (4)
    Adaptive smoothing method (ASM) [43]: interpolates and smooths the values using two different smoothing kernels: one for free-flowing traffic and the other for congested traffic. The results are then combined using a $\tanh$ activation function.
  5. (5)
@@ 5.3 Code and tools
The model is coded using PyTorch, leveraging CUDA for accelerated computation. All experiments are performed on a machine with two NVIDIA RTX A6000 GPUs. The code is open-sourced and available at i24motion.org, facilitating future research and enabling other researchers to build upon our work.
@@ 5.4 Evaluation Metrics
  1. (1)
    Wasserstein distance (WD) [66]: The Wasserstein distance quantifies the difference between two probability distributions. A smaller distance suggests that the distribution is more closely aligned with the reference distribution.
  2. (2)
    Root mean squared error (RMSE): in miles per hour
  3. (3)
    Mean absolute percentage error (MAPE): as a percentage of the true speed
Table 3: Evaluation matrices on different data refinement methods and our improvements over the second best method
Metric Average Nearest Linear ASM Ours Improvement
Training
WD 20.74 1.92 2.82 3.08 1.24 35.42%
RMSE 23.19 9.36 7.87 7.38 5.38 27.10%
MAPE 1.37 0.47 0.48 0.41 0.16 60.98%
Validation
WD 20.35 1.97 2.86 3.12 1.38 29.95%
RMSE 22.77 9.56 8.11 8.05 6.81 15.40%
MAPE 1.33 0.49 0.49 0.43 0.28 29.95%
Testing
WD 19.39 1.81 2.67 3.18 1.59 12.15%
RMSE 21.98 9.88 8.31 7.87 7.17 8.89%
MAPE 1.41 0.54 0.54 0.48 0.32 33.33%
6 Results and discussions
@@ 6.1 Overall performance
Table 3 presents a detailed evaluation comparing the performance of our method with various baselines. The results clearly indicate that our method outperforms the baseline methods across all evaluation metrics and datasets. Notably, the superior performance of our method in minimizing the Wasserstein Distance (WD), 35.42% in training, 29.95% in validation and 12.15% in testing over the second-best methods, underscores its efficacy in accurately reconstructing speed distributions, which is crucial for stop-and-go wave reconstruction. The substantial improvement in the WD metric suggests that our method captures the correct patterns, even if there are minor spatial misplacement, which is an acceptable limitation as confirmed by the travel time estimates.
In terms of RMSE, our method demonstrates the improvements, with a 27.10% advantage in training, 15.40% in validation, and 8.89% in testing, indicating its robustness in producing refined data with minimal error. The MAPE metric reveals even more pronounced benefits, with our method achieving a 60.98% advantage in training, 34.88% in validation, and 33.33% in testing. These results highlight the overall efficacy of our approach in improving traffic reconstruction across various metrics.
@@ 6.2 Congested traffic reconstruction
Figure 7: Distribution of accumulated speed profiles: ours (top), reference (middle), ASM (bottom)
@@ 6.4 Travel time reconstruction
By utilizing the generated speed field, travel time can be estimated through the use of virtual vehicles [67; 28]. For each scenario, a number of vehicles are introduced into the speed field to create virtual trajectories, allowing for the estimation of travel time through these virtual vehicles. In our case, virtual vehicles are deployed into the speed field every 10 seconds, with the corresponding virtual trajectories updated at a frequency of 1Hz.
@@
18-30 mph (37 total)
mean (seconds) 581.1 509.7 603.9
relative error - -12% +4%
@@
mean (seconds) 383.9 358.5 387.1
relative error - -7% +1%
@@
For the 18-30 mph speed range, which includes a total of 37 samples, the mean travel time estimated by our model is 603.9 seconds with a standard deviation of 89.0 seconds. Compared to the reference mean travel time of 581.1 seconds, our model shows a relative error of +4%, whereas the ASM model exhibits a relative error of -12%. This indicates that while our model slightly overestimates the travel time in this speed range, it does so with reasonable accuracy. In the 30-50 mph speed range, consisting of 25 samples, our model estimates a mean travel time of 387.1 seconds with a standard deviation of 62.5 seconds. The reference mean travel time is 383.9 seconds, resulting in a relative error of +1% for our model. The ASM model, on the other hand, has a relative error of -7%. This suggests that our model performs with high accuracy in this speed range, with minimal deviation from the reference values. The findings suggest that our model is robust across various speed ranges, especially good in dealing with the challenging stop-and-go traffic conditions.
@@ 6.5 Wave speed estimation
We apply the wave analysis tools developed in [24] to validate the wave speed for wave front and tails. Wave fronts are identified as the spatio-temporal points where a virtual vehicle speed drops to 16 mph, while wave tails are defined as the points where the speed increases back above 16 mph.
@@
mean speed (mph) 11.97 13.28 13.31
speed st.dev. (mph) 2.10 2.22 2.78
@@
mean speed (mph) 11.51 12.79 12.48
speed st.dev. (mph) 1.89 2.16 2.69
As can be seen from the Table 5, our reconstructed wave speed is very close to the ASM, which is predefined as 13 mph when preprocessing the low spatio-temporal resolution data. A likely cause is the conditional diffusion process on low spatio-temporal resolution data with preprocessing, particularly the upsampling using ASM, which appears to dominate the observed phenomena. This represents a limitation of the current work, necessitating further experiments to explore and mitigate the impact of preprocessing techniques on wave speed estimation.
@@ 6.6 Events and crashes
We conducted a manual inspection of all outputs from the validation dataset to identify any apparent errors produced by our models. Figure 10 shows 3 of the typical failed examples in the validation datasets. The initial two rows of the examples indicate that there are crashes/events concealed within the stop-and-go waves. These are not detectable by conventional sensors and, consequently, cannot be retrieved using our methods, which is acceptable. Future research could explore the potential of incorporating event context information into the model as a form of prompt for generation. The final row of the example indicates that data is missing for half of the observation period. The interpolation is degraded by backfilling interpolation, which results in incorrect input and thus inaccurate output. Future research could also investigate whether the generative framework can be useful for data imputation.
@@ 7 Conclusion
Our proposed approach is validated using a new dataset I24-WaveX which provides a comprehensive ground for testing and comparing our model against various baselines. The results clearly show that our method outperforms the existing techniques across several evaluation metrics, notably in minimizing the Wasserstein Distance (WD) by 29.95% on validation dataset and improving Root Mean Square Error (RMSE) and Mean Absolute Percentage Error (MAPE). Despite spatio-temporal mismatches, our method significantly advances beyond current baseline methods on stop-and-go wave reconstructions by accurately capturing the wave patterns, leading to better estimations of travel times and stop-and-go wave speeds.
However, limitations are also identified, particularly in scenarios involving crashes and complex traffic events, where fine-grained details do not align perfectly with high-fidelity data.
In conclusion, our research contributes to the field of traffic data enhancement by providing a viable method to improve the resolution of conventional traffic sensors. The open-sourcing of our data, trained model and code can also facilitate further research and practical applications with generative AI. Future work will focus on improving our model to address its current limitations, ensuring greater accuracy and reliability in more diverse traffic states.
Acknowledgements
This work was supported by the National Science Foundation (NSF) under Grant No. 2135579 (Work, Sprinkle), and 2111688 (Sprinkle), the U.S. Department of Transportation under Grant No. 693JJ22140000Z44ATNREG3202, and the Tennessee Department of Transportation under Grant No. RES2023-20. It is also supported by the USDOT Dwight D. Eisenhower Fellowship program under Grant Agreement 693JJ32445065 (Richardson). We acknowledge the Tennessee Department of Transportation (TDOT) for providing the data used in this research. We are grateful to Dr. Yanbing Wang from Argonne National Laboratory for her contributions to the I-24 MOTION project. We also appreciate the insights for spatio-temporal modeling from Austin Coursey and Prof. Tyler Derr at Vanderbilt University. The views expressed herein do not necessarily reflect those of the U.S. Department of Transportation, the Tennessee Department of Transportation, or the United States Government.

### v2

@@ 4.1 I-24 MOTION data
Trajectory data produced by I-24 MOTION was used to create fine-grained mean speed field data for this work. A selection of 33 days were selected, recorded from May 17 to July 17, 2024; for each day the morning rush hour period (6:00 or 5:30 AM to 10:00 AM) was considered as this period has the most dynamic and variable traffic conditions. Figure 3 shows the raw trajectory data for each selected day. The selected trajectory data was then sampled to create a fine-grained mean speed field with 4 second and 0.02 mile bin resolution. For cells with no data and cells with inconsistent values, speed was interpolated using an adaptive smoothing method [46; 29; 66], with parameters shown in Table 2.
@@
Table 1 provides an overview of the key metrics in the I-24 MOTION dataset, including Vehicle Miles Traveled (VMT), Vehicle Hours Traveled (VHT), and the mean speed (MS, the vehicle-miles-weighted speed) [67]. It also outlines the strategy for splitting the data into training and validation sets, with a 75% allocation for training and 25% for validation. Note that the testing set comprises data from 5 days with distinctive congestion patterns due to observable events, which are excluded from both the training and validation sets to avoid model training imbalance. The validation set is used during training to tune hyperparameters and monitor the performance, providing feedback to prevent overfitting (e.g. early stopping). The test set is kept completely separate from the training process and is used only to evaluate the model generalization on unseen data.
@@ 4.2 I-24 RDS data
The I-24 RDS data is obtained from the I-24 radar detection system (RDS) [68; 69]. The system collects lane-level local traffic measurements including speed, occupancy, and volume, aggregated at 30-second intervals. All the radar sensors utilized in this study were meticulously calibrated. For the purposes of this study, we specifically selected the subset of sensors that align spatially with the I-24 MOTION testbed as shown in Figure 4. A total of 11 RDS sensors are aligned in this section. Two additional sensors located at both ends are selected to help with reconstruction.
@@ 4.3 Dataset processing
$\sigma$(mile) smoothing width in time coordinate 0.02 0.10
$\tau$(second) smoothing width in space coordinate 4 9
$c_{\text{free}}$(mph) wave speed in free traffic -12 -12
$c_{\text{cong}}$(mph) wave speed in congested traffic 45 45
$V_{\text{thr}}$(mph) crossover from congested to free traffic 40 40
$\Delta V$(mph) transition width between congested and free traffic 10 10
Dataset augmentation. To augment the dataset, we employed a sliding window strategy, generating multiple overlapping sub-datasets from the original daily dataset matrix. Starting with an original tensor of size 3600 $\times$ 200, we created multiple 200 $\times$ 200 sub-tensors. A 200 $\times$ 200 window slides across the larger matrix with a defined step size, producing a new 200 $\times$ 200 tensor at each step. In this study, we used a sliding window step size of 10, corresponding to 40 seconds in time. For the evaluation part, the augmentation via sliding window is not used.
@@ 4.4 Data Availability
The dataset associated with this paper, WaveX, would be a new release compared to our previous INCEPTION dataset. All days of high-resolution mean speed field data, as well as all corresponding days of RDS data, will be made available at i24motion.org.
@@ 5.1 Task specific architectural details
In our problem, the coarse-grained data $\mathbf{r}$ is preprocessed with upsampling to be $\mathbf{\tilde{r}}$ (as shown in Figure 5). The upsampled data is then concatenated with the noisy fine-grained speed data $\mathbf{m}_{t}$. This process generates a 4-channel tensor as the model input, while the model output is a single-channel tensor $\mathbf{m}_{t-1}$. The UNet [64] architecture employed in this paper comprises an initial level with 64 channels, followed by levels with 128, 256, and 512 channels, culminating in a bottleneck level with 1024 channels, as shown in Figure 5. Each block within this architecture consists of three convolutional ResNet blocks augmented with a multi-head attention mechanism, leading to a total of 130 million trainable parameters.
@@
In this paper, we define a sequence of $\gamma$ values, which are essential for our diffusion process. These values are uniformly distributed with $p(\gamma)$ within a specified range for $T$ steps, starting from $\gamma_{\text{start}}=0.001$ to $\gamma_{\text{end}}=0.02$. The total number of values in this sequence is equal to the number of time steps, $T=500$, in our diffusion process. The learning rate for the optimizer of the neural denoising model is set at $2\times 10^{-4}$. The velocity $v_{c}$ is set to 15 mph, and the parameter $\omega$ is set to 0.8. An early stopping strategy is employed when the validation loss begins to increase. The best performance on the validation dataset is retained for inference purposes.
@@ 5.2 Benchmarked methods
Figure 6: Comparison of baseline methods and our reconstruction against the ground truth.
  1. (i)
    Average: All data points are computed as the average of the available samples, under the assumption that the data is uniformly distributed across both space and time.
  2. (ii)
@@
  4. (iv)
    Adaptive smoothing method (ASM) [46; 49]: interpolates and smooths the values using two different smoothing kernels: one for free-flowing traffic and the other for congested traffic. The results are then combined using a $\tanh$ activation function.
  5. (v)
@@
A visual demonstration of each baseline method, alongside our reconstruction and the ground truth, is presented in Figure 6.
@@ 5.3 Code and tools
The model is coded using PyTorch [70], leveraging CUDA for accelerated computation. All experiments are performed on a machine with two NVIDIA RTX A6000 GPUs and AMD Ryzen Threadripper 3960X 24-Core Processor. The code will be open-sourced , facilitating future research and enabling other researchers to build upon our work.
Table 3: Evaluation matrices on baselines highlighting our improvements over the second-best approach. In the results, the second-best method is indicated with underline, while the best-performing method is marked in bold.
Metric Average Nearest Linear ASM Ours Improvement
Training
WD 20.88 1.90 2.72 2.36 1.05 47.22%
RMSE 23.40 9.43 8.01 6.89 4.44 35.49%
MAPE 1.50 0.48 0.49 0.37 0.14 62.57%
Validation
WD 20.51 1.94 2.75 2.36 1.28 34.02%
RMSE 23.00 9.48 8.09 6.97 5.94 14.81%
MAPE 1.45 0.48 0.48 0.37 0.25 32.18%
Testing
WD 19.68 1.79 2.59 2.27 1.48 17.32%
RMSE 22.33 9.82 8.35 7.22 6.25 13.40%
MAPE 1.55 0.53 0.54 0.43 0.28 34.87%
@@ 5.4 Evaluation Metrics
  1. (i)
    Wasserstein distance (WD) [71]: The Wasserstein distance quantifies the difference between two probability distributions. A smaller distance indicates a closer alignment with the reference distribution.
  2. (ii)
    Root Mean Squared Error (RMSE): Unit in miles per hour (mph), RMSE measures the average magnitude of the errors between reconstructed and actual values, with greater weight assigned to larger errors. Lower RMSE values indicate more accurate reconstructions.
  3. (iii)
    Mean Absolute Percentage Error (MAPE): Reported as a percentage, MAPE reflects the average absolute difference between reconstructed and actual speed relative to the actual speed. Lower MAPE values indicate more accurate reconstructions.
6 Results
@@ 6.1 Overall performance
Table 3 presents a detailed evaluation comparing the performance of our method with various baselines. The results clearly indicate that our method outperforms the baseline methods across all evaluation metrics and datasets. Notably, the superior performance of our method in minimizing the Wasserstein Distance (WD), 47.22% in training, 34.02% in validation and 17.32% in testing over the second-best methods, underscores its efficacy in accurately reconstructing speed distributions, which is crucial for stop-and-go wave reconstruction. The substantial improvement in the WD metric suggests that our method captures the correct patterns, even if there are minor spatial misplacement, which is an acceptable limitation as confirmed by the travel time estimates.
In terms of RMSE, our method demonstrates the improvements, with a 35.49% advantage in training, 14.81% in validation, and 13.40% in testing, indicating its robustness in producing refined data with minimal error. The MAPE metric reveals even more pronounced benefits, with our method achieving a 62.57% advantage in training, 32.18% in validation, and 34.87% in testing. These results highlight the overall efficacy of our approach in improving traffic reconstruction across various metrics.
@@ 6.2 Congested traffic reconstruction
Figure 8: Distribution of accumulated speed profiles on the validation dataset: ours (top), reference (middle), ASM (bottom)
@@ 6.4 Travel time reconstruction
By utilizing the generated speed field, travel time can be estimated through the use of virtual vehicles [72; 29]. For each scenario, a number of vehicles are introduced into the speed field to create virtual trajectories, allowing for the estimation of travel time through these virtual vehicles. In our case, virtual vehicles are deployed into the speed field every 10 seconds, with the corresponding virtual trajectories updated at a frequency of 1Hz.
@@
18-30 mph (36 total)
mean (seconds) 579.9 512.1 596.7
relative error - -12% +3%
@@
mean (seconds) 385.0 360.5 390.5
relative error - -6% +1%
@@
For the 18-30 mph speed range, which includes a total of 36 samples, the mean travel time estimated by our model is 596.7 seconds with a standard deviation of 81.3 seconds. Compared to the reference mean travel time of 579.9 seconds, our model shows a relative error of +3%, whereas the ASM model exhibits a relative error of -12%. This indicates that while our model slightly overestimates the travel time in this speed range, it does so with reasonable accuracy. In the 30-50 mph speed range, consisting of 25 samples, our model estimates a mean travel time of 390.5 seconds with a standard deviation of 63.8 seconds. The reference mean travel time is 385.0 seconds, resulting in a relative error of +1% for our model. The ASM model, on the other hand, has a relative error of -6%. This suggests that our model performs with high accuracy in this speed range, with minimal deviation from the reference values. The findings suggest that our model is robust across various speed ranges, especially good in dealing with the challenging stop-and-go traffic conditions.
@@ 6.5 Wave speed estimation
We apply the wave analysis tools developed in [25; 66] to validate the wave speed for wave front and tails. Wave fronts are identified as the spatio-temporal points where a virtual vehicle speed drops to 15 mph, while wave tails are defined as the points where the speed increases back above 15 mph.
@@
mean speed (mph) 12.01 12.22 12.69
speed st.dev. (mph) 2.14 1.82 2.49
@@
mean speed (mph) 11.24 12.16 11.81
speed st.dev. (mph) 1.65 1.90 1.88
As can be seen from the Table 5, our reconstructed wave speed is very close to the ASM, which is predefined as 12 mph when preprocessing the low spatio-temporal resolution data. Our reconstruction captures the varying speeds of the wave fronts and wave tails, highlighting its effectiveness in reconstructing the evolving patterns of stop-and-go waves. In contrast, the ASM produces relatively consistent speeds for both the wave front and tail, reflecting its predefined assumptions rather than the dynamic nature of real traffic patterns. However, there is a mismatch in the estimated mean speed between the wave fronts and wave tails. Addressing this mismatch may require further development, potentially by incorporating mechanisms to explicitly enforce wave speed consistency, such as integrating relevant constraints into the loss function.
7 Discussions
@@ 7.1 Events and crashes
We conducted a manual inspection of all outputs from the validation dataset to identify any apparent errors produced by our models. Figure 11 shows 3 of the typical failed examples in the validation datasets. The initial two rows of the examples indicate that there are crashes/events concealed within the stop-and-go waves. These are not detectable by conventional sensors and, consequently, cannot be retrieved using our methods, which is acceptable. Future research could explore the potential of incorporating event context information into the model as a form of prompt for generation. The final row of the example indicates that data is missing for half of the observation period. The ASM interpolation leverages the non-compact kernel to extend the reconstructed waves; however, this leads to an overestimation of wave propagation, which in turn causes our method to reconstruct the wave in the leftmost section for a longer duration than observed.
@@
7.2 Applicability and transferability
The learned mapping from coarse-grained sensor data to fine-grained traffic patterns can be employed to reconstruct stop-and-go waves on other freeway segments, provided the following prerequisites are met:
  1. (i)
    The sensor network need have a similar spatial resolution and employ a 30-second temporal aggregation for sensors that measure speed, occupancy, and flow.
  2. (ii)
    The sensors need be properly configured to ensure data compatibility with the learned mapping.
We demonstrate the trained model perform on an untrained segment in Figure 13 on another road stretch of I-24 (Mile Marker 54.5 to Mile Marker 58.5).
Figure 13: Transferability of the model to another freeway stretch (Mile Marker 54.5 to Mile Marker 58.5) on I-24 demonstrated on June 11, 2024. The model is trained on the I-24 MOTION stretch (Mile Marker 58.7 to Mile Marker 62.7).
The proposed training process is transferable, and the model developed in this study can serve as an effective initialization for re-training on a different freeway segment, enabling fine-tuning [73] for new scenarios and different geographic locations.
7.3 Challenges on data quality
The quality of the data, particularly the accuracy of the ground truth fine-grained data, plays a critical role in model performance. While sporadic issues such as communication loss or data anomalies can often be mitigated through preprocessing in our upsampling module (ASM), systematic biases present a more persistent challenge. In the following, we provide a detailed discussion and evidence on the risks and impacts of data bias.
  1. (i)
    Biased training data, unbiased or biased input: It would clearly lead to biased generated outputs. This issue has been discussed in the generative modeling literature [74], where the impact of training data bias on model generation distribution have been investigated [75; 76].
  2. (ii)
    Unbiased training data, biased input: In our experiments, we introduce bias into the input data by adding Gaussian noise to the coarse-grained observations. Specifically, the noise is drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, where the mean $\mu$ and standard deviation $\sigma$ are independently sampled from the discrete set $\{0,1,2,3,4,5\}$. This formulation allows us to evaluate the model performance on varying degrees and types of input bias.
    Figure 14: RMSE degradation under varying levels of Gaussian noise added to the raw data.
    Figure 14 demonstrates the degradation in model performance under varying levels of Gaussian noise, evaluated using data from June 3, 2024. The results indicate that the quality of input data impacts model performance. Specifically, bias in the mean value of the added noise (value bias) tends to have a more pronounced impact on RMSE degradation compared to random bias introduced through variations in the standard deviation. This suggests that systematic shifts in the data distribution are more detrimental to the model than random fluctuations of comparable scale. The overall RMSE degradation remains bounded within 30% across the tested noise configurations.
    For sporadic issues such as communication loss or data anomalies, the upsampling component ASM is specifically designed to address these challenges, as illustrated in the bottom example of Figure 7 and supported by prior studies [48; 49].
@@ 8 Conclusion
Our proposed approach is validated using a new dataset WaveX which provides a comprehensive ground for testing and comparing our model against various baselines. The results clearly show that our method outperforms the existing techniques across several evaluation metrics, notably in minimizing the Wasserstein Distance (WD) by 34.02% on validation dataset and improving Root Mean Square Error (RMSE) and Mean Absolute Percentage Error (MAPE). Despite spatio-temporal mismatches, our method significantly advances beyond current baseline methods on stop-and-go wave reconstructions by accurately capturing the wave patterns, leading to better estimations of travel times and stop-and-go wave speeds. However, limitations are also identified, particularly in scenarios involving crashes and complex traffic events, where fine-grained details do not align perfectly with high-fidelity data.
In conclusion, our research contributes to the field of traffic data refinement by providing a viable method to improve the resolution of conventional traffic sensors. The open-sourcing of our dataset, trained model, and code enable further research and practical applications of generative AI in freeway operations.
Acknowledgments
We are grateful for the thoughtful insights provided by our anonymous reviewers. This work was supported by the National Science Foundation (NSF) under Grant No. 2135579 (Work, Sprinkle), and 2111688 (Sprinkle) and the Tennessee Department of Transportation under Grant No. RES2023-20 and Grant No. OTH2023-01F-01. It is also supported by the U.S. Department of Transportation Dwight D. Eisenhower Fellowship program under Grant Agreement 693JJ32445065 (Richardson). We acknowledge the Tennessee Department of Transportation (TDOT) for providing the data used in this research. We are grateful to Prof. Yanbing Wang from Arizona State University for her contributions to the I-24 MOTION project. We also appreciate the insights for spatio-temporal modeling from Austin Coursey and Prof. Tyler Derr at Vanderbilt University. The views expressed herein do not necessarily reflect those of the U.S. Department of Transportation, the Tennessee Department of Transportation, or the United States Government.